# A comparative study of different TSO-DSO coordination in the reserve market

### v1

@@ III-B TSO model
TSO leverages the DER flexibility to minimize the cost they pay to balance the system per each ISP. TSO objective function is described in Eq. 18 and the constraints following need to be satisfied:
$F_{t}^{TSO}=\sum_{m\in M}\ E_{m,t}^{a,\uparrow}\lambda_{m}^{a,\uparrow}+\sum_{k\in K}E_{k,t}^{a,\downarrow}\lambda_{k}^{a,\downarrow}$ $+E_{t}^{r,\uparrow}\lambda_{t}^{\uparrow}-E_{t}^{r,\downarrow}\lambda_{t}^{\downarrow}$ (18)
@@
In the offering process, TSO defines a list of selected bids, with prices ranging from low to high until the required quantity is reached to ensure balance. Therefore, the bid prices ($\lambda_{m}^{a,\uparrow}$, $\lambda_{k}^{a,\downarrow}$) for regulating upward and for regulating downward by aggregator $m$ and $k$ are factored into the cost function of TSO. Besides, $\lambda_{t}^{\uparrow}$ and $\lambda_{t}^{\downarrow}$ are prices of reserve capacity for regulating upward and for regulating downward from reserve units, which is assumed equal the balancing prices (imbalance price is calculated by the balancing price and most of the time when only upward or downward regulation in an ISP, these prices are identical). The variables $E_{m,t}^{a,\uparrow}$,$E_{k,t}^{a,\downarrow}$ are the volume of flexibility provided by aggregator $m$ and aggregator $k$ at step $t$. $E_{t}^{reg,\uparrow}$ and $E_{t}^{reg,\downarrow}$ are the total volume activated for regulation and $E_{t}^{r}$ is the volume of reserve capacity uses. $M$ and $K$ are the number of aggregator flexibility units.
Equations 19 and 20 ensure that the total volume provided by the DER flexibility and reserve units is equal to the volume required for upward or downward regulation by the TSO. Equations 21 and 22 limit the volume activated by the TSO within the boundary that the aggregator can provide. Equations 23 and 24 represent the sign that the volume regulation for upward regulation is greater than zero and the volume for downward regulation less than zero. While equations 25, 26, and 27 are the boundaries calculated from the output of aggregator models.
@@ III-C DSO model
The DSO will receive the MOL from the TSO or the boundaries from the aggregator. Along with their demand forecast, DSO will perform the power flow analysis to check if any congestion occurs in their network. The DSO’s objective function is to minimize the potential cost they have to pay to resolve congestion per ISP which is shown in Eq. 28:
$F^{DSO}_{t}=\sum_{x\in\Omega_{x}}(V_{x,t}^{\uparrow}C_{t}^{\uparrow})-(V_{x,t}^{\downarrow}C_{t}^{\downarrow})$ (28)
In this case, only congestion caused by the activation of downward or upward regulation from the EVs aggregators within the management of DSO is considered. Therefore, flexibility used to solve congestion is from the same aggregation unit as TSO. $V_{x,t}^{\uparrow},V_{x,t}^{\downarrow}$ are the volume of flexibility used for managing congestion at node $x$, which is in a set of buses $\Omega_{x}$. $M$ and $K$ is a subset of $\Omega_{x}$. $C_{t}^{\uparrow}$ and $C_{t}^{\downarrow}$ are congestion prices of flexibility for adjusting upward and downward at step $t$.
In the TSO-DSO hybrid-managed model, the DSO uses Eq. 29 and Eq. 30 to update the demand $P_{i,t}^{d^{\prime}}$ and generation $P_{i,t}^{g^{\prime}}$ in the next two ISPs with the volumes used by the TSO from the MOL $E_{x,t}^{a,\uparrow},E_{x,t}^{a,\downarrow}$. Then, DSO performs the optimal power flow calculation based on Eq. 31-38 to determine the flexibility volume $V_{x,t}^{\uparrow},V_{x,t}^{\downarrow}$ to resolve congestion. If the optimization calculation gives a feasible solution, meaning that the amount of flexible resources from the aggregators is sufficient to resolve the congestion caused by the use of the TSO’s service. Otherwise, the boundary for the TSO is divided by $step$-th in 39 and 40. TSO can only use resources within this new limitation. The process is then performed until the DSO finds the optimal solution or reaches the stopping condition after five divisions, the $E_{m,t}^{a,\uparrow^{\prime}},E_{k,t}^{a,\downarrow^{\prime}}$ are reset to 0.
@@
Differently, in the DSO-managed coordination scheme, DSO updates the demand and generation at each node with $\overline{E_{x,t}^{a,\uparrow}},\underline{E_{x,t}^{a,\downarrow}}$ from the aggregator.
Active power balance at node $x$ at step $t$ is presented in Eq. 31, in which, $J_{x}$ is the subset of nodes that are connected to node $x$. This constraint needs to be satisfied for all nodes in the network. Similar to TSO model, the constraints on flexibility boundaries are presented in Eq. 32, Eq. 33 and Eq. 34. Power flow symmetry condition in Eq. 35 means that the flow of power from node $x$ to node $y$ ($Flow_{xy}$) is opposite to the value of flow from node $y$ to node $x$. This value is determined by the base power for the network (in MVA) $Base_{MVA}$, the susceptance of line in per unit $B_{xy}$ and the voltage phase angle difference $\theta_{x,t}$ and $\theta_{y,t}$. $B_{xy}$ is calculated by resistance and reactance of the line in Eq. 38. Moreover, the rated value of each branch ($Rated_{xy}$) in Eq. 36 multiplied with the power factor $PF$ forms the thermal rating threshold for each branch. The power factor $PF$ is considered 98% in this study. The above problems are mixed-integer linear programming (MILP) problems and are solved by the Gurobi solver in Pyomo.
@@ IV Case studies
The two proposed coordinations with DERs’ flexibility provided by the aggregator are validated using the real distribution network in the Bleiswijk station in the northern part of the Netherlands. This network includes 4 step-down two-winding transformers from 25kV to 10.5kV, 182 busbars, 867 terminals, 224 lines, 25 synchronous machines, 170 substations, and 166 loads. The total installed capacity is 66.06 MW with a total spinning reserve is 11.36 MW. The PV penetration rate is 12.85% ( 8.5MW) from 16 installation points. The aggregator for upward is connected to nodes 12, 42, 145, 146, and 147 and for downward is connected to nodes 15, 18, 41, 179, and 183 in the network. Each aggregator manages 100 EVs and the aggregated flexibilities are identical. The only difference is their bid prices, which are presented in Tables I. One week in summer and one week in winter, different BRP prices are simulated to consider the aggregator behavior in price fluctuations and seasonal changes.
TABLE I: The DER flexibility for upward regulation
@@ IV-A Aggregator behavior by market price and season
One week in summer, 2023 from June 26th to July 2nd and one week in winter, 2024 from February 19th to February 25th is selected. With a BRP price of 30 €/MW and a consumer price of 85 €/MW, respectively, which is lower and higher than the average day-ahead price this week. Both in winter and summer, the volume activated for upward and downward is concentrated when the DA price is high, the DA volume is concentrated when the DA price is low, which is presented in Fig. 3 and Fig. 4, respectively. The occurrence of downward service is very low and has the lowest total volume, followed by upward service, and the highest is the volume purchased from the DA market. Besides, the downward volume is densely distributed in the region where the upward price is zero and vice versa sparsely distributed in the high upward price region. The same behavior is for the distribution of upward volume versus downward price. In most cases in reality, when the system is in deficit, the upward regulation price will increase to attract more volume from service providers, the downward regulation at that time will aggravate the situation of the system. Conversely, there is no need for increased volume when the system is surplus, and the downward regulation price goes very low. The distribution of buy volume by DA price and downward volume by downward price in summer is wider than in winter due to energy surplus from solar power plants with zero-marginal prices leading to many negative price cases for DA and downward price. On the opposite, it’s wider for upward price range caused by energy shortage in winter.
Figure 3: Distribution of volume by price in winter.
Figure 4: Distribution of volume by price in summer.
From the pie chart in Fig. 5 show that the total volume provided for upward service (blue part) in summer is lower than in winter due to the need for more energy in winter. Downward volume (green part) drops sharply in winter when BRP prices are high. The total volume for providing service is higher when the BRP price is small, especially for upward service in winter. Moreover, aggregator profit is inversely proportional to BRP price. The majority of the aggregator’s profits come from providing upward services leading to the aggregator’s profit in winter being higher than in summer.
Figure 5: Distribution of volume by season with different BRP price.
@@
BRP price impact is clearly illustrated in the total volume upward, downward regulation, and buy from DA market of each EV. Fig. 6 and Fig. 7 show the difference in the total volume downward when the BRP price is 200 €/MWh and 30 €/MWh. The orange bar presents the downward service almost disappears. Besides, the volume for upward (blue bar) is also decreased, leading to the volume bought from the DA market (green bar) being reduced, because they don’t have to buy more to compensate for the volume discharge when providing upward service.
Figure 6: Volume by type of EV1 with BRP price = 30€/MWh.
Figure 7: Volume by type of EV1 with BRP price = 200€/MWh.
IV-B Grid constraint evaluation
@@
Providing services brings more benefits to aggregators and prosumers. However, a problem arises when the TSO activates the service without considering the distribution network constraints, and multiple aggregators react at the same time, resulting in congestion within the distribution network. The green line in Fig. 8 shows that activating the TSO balancing service at ISP 44 causes overloading in the main line straight to the connection point with the external grid. The congestion above is due to the regulation upward from TSO, the line loading is nearly 100% because of the upward regulation when the distribution grid is in a low-demand situation, causing a reversed flow from the distribution grid to the transmission grid. This will cause rapid deterioration of the device or damage if the overload condition is prolonged. The dashed red line presents the limitation of the line loading, which is 95% in this study. With the coordination between TSO and DSO, after the DSO updates the MOL sent by TSO, they limit the volume of flexibility that TSO can use. Loading after TSO use flexibility with the new boundary is now below the limit constraints, which are presented by the orange line.
@@
In the DSO-managed model, DSO based on the demand forecast for the next 2 ISP, performs power flow analysis with all the flexibility volume sent by the aggregator. The result is presented in the green line in Fig. 9. Similar to the previous case, there is overloading occurs in the ISP 44. Then DSO has to update the boundary and send the new boundary to TSO. The loading of the measuring point is now the orange line, which shows that the limit constraint is satisfied.
However, the different between two cases is that the volume of flexibility that the aggregator can provide in the second case after DSO updates the boundary is less than the previous one, due to DSO not knowing exactly where TSO will call service, they will gradually reduce the boundary for all flexibility units until the congestion can be resolved by the available flexibility. This leads to the loss of profit for aggregators if they reserve energy for providing service in real-time, but this volume is not valid to provide anymore.
@@ V Conclusion
To effectively harness the potential of the DERs services in the current context, coordination between TSOs and DSOs is necessary to maximize social welfare as well as minimize the risks of conflicting use cases. Notably, DSOs play a more important role in validating services before being utilized by TSOs. This validation ensures that grid constraints are taken into account when providing services from aggregators under the DSO’s management. Thereby, mitigating the potential risks of unexpected overload caused by the simultaneous reaction of aggregators with prices and requests from TSOs. Aggregator’s objective is to maximize their profits by optimizing bidding strategies. Analyzing the service provision from aggregator shows that similar behavior from multiple aggregators in neighboring regions in response to attractive very high upward regulation price or very low downward regulation price will lead to excessive reversed power flow from their connection points to the higher voltage grid or vice versa. Besides, the price the aggregator contracts with BRP for their real-time deviation has a great impact on the decision to provide services and the final profit. The EVs aggregator will gain more profit from providing the upward regulation service than downward and therefore it is clear that winter will be more profitable due to the higher energy demand and frequency of upward service calls. The coordination between TSO and DSO in the offering process allows the provision of services within the grid constraints. When the TSO has the priority in receiving information in the hybrid-managed model, the DSO updates the boundaries based on the TSO’s selected bids. Conversely, in the DSO-managed model, unused bids are considered when DSOs resolve congestion resulting in a larger portion of services being excluded in this scenario compared with the first one. Therefore, the hybrid-managed gives an advantage with more volume provided and more benefit to the aggregator than the DSO-managed coordination scheme. The risks associated with inaccurate forecasts of demand, price, and EV sessions significantly influence the decisions of each participant. These challenges will be addressed in future studies.

### v2

@@ III-B TSO model
TSO leverages the DER flexibility to minimize the cost of balancing the system for each ISP by the function below:
$min\sum_{m\in M}\ E_{m,t}^{a,\uparrow}\lambda_{m}^{a,\uparrow}-\sum_{k\in K}E_{k,t}^{a,\downarrow}\lambda_{k}^{a,\downarrow}$ $+E_{t}^{r,\uparrow}\lambda_{t}^{\uparrow}-E_{t}^{r,\downarrow}\lambda_{t}^{\downarrow}$ (11)
The TSO has to satisfy the following constraints:
@@
In the offering process, TSO compiles a list of selected bids, ordered by price from low to high, until the required quantity is reached to ensure balance. Therefore, the bid prices ($\lambda_{m}^{a,\uparrow}$, $\lambda_{k}^{a,\downarrow}$) for regulating upward and for regulating downward by aggregator $m$ and $k$ are factored into the cost function of TSO. Besides, $\lambda_{t}^{\uparrow}$ and $\lambda_{t}^{\downarrow}$ are prices of reserve capacity for upward and downward regulation from reserve units. These prices are assumed to be equal to the balancing prices (the imbalance price is derived from the balancing price, and in most cases, when only upward or downward regulation is required, these prices are identical). The variables $E_{m,t}^{a,\uparrow}$,$E_{k,t}^{a,\downarrow}$ are the volume of flexibility provided by aggregator $m$ and aggregator $k$ at step $t$. $E_{t}^{reg,\uparrow}$ and $E_{t}^{reg,\downarrow}$ are the total volume activated for regulation and $E_{t}^{r}$ is the volume of reserve capacity used. $M$ and $K$ are the number of aggregator flexibility units.
Eq. 12 and Eq. 13 ensure that the sum of the volume provided by the DER flexibility and reserve units is equal to the volume required for upward or downward regulation by the TSO. Equations 14 and 15 limit the volume activated by the TSO within the boundary that the aggregator can provide. Equations 16 and 17 represent the sign that the volume regulation for upward regulation is greater than zero and the volume for downward regulation is less than zero. Eq. 18, Eq. 19, and Eq. 20 present the calculation of the boundaries from the output of aggregator models.
@@ III-C DSO model
The DSO will receive the MOL from the TSO or the boundaries from the aggregator. Along with their demand forecast, DSO will perform the power flow analysis to check if any congestion occurs in the network. The DSO’s objective function is to minimize the potential cost they have to pay to resolve congestion per ISP, which is shown in Eq. 21:
$min\sum_{x\in\Omega_{x}}(V_{x,t}^{\uparrow}C_{x}^{\uparrow})-(V_{x,t}^{\downarrow}C_{x}^{\downarrow})$ (21)
In this case, congestion is assumed to arise solely from the activation of downward or upward regulation from the EV aggregators. Consequently, the flexibility deployed to mitigate congestion is drawn from the same aggregation unit as TSO. $V_{x,t}^{\uparrow},V_{x,t}^{\downarrow}$ are the volume of flexibility used for managing congestion at node $x$, which is in a set of buses $\Omega_{x}$. $M$ and $K$ is a subset of $\Omega_{x}$. $C_{x}^{\uparrow}$ and $C_{x}^{\downarrow}$ are congestion prices of flexibility for adjusting upward and downward at node $x$, this price is assumed to be equal to the price for the reserve bid by using the same flexibility unit.
In the TSO-DSO hybrid-managed model, the DSO uses Eq. 22 and Eq. 23 to update the demand $P_{x,t}^{d^{\prime}}$ and generation $P_{x,t}^{g^{\prime}}$ in the next two ISPs with the volumes used by the TSO from the MOL $E_{x,t}^{a,\uparrow},E_{x,t}^{a,\downarrow}$. Then, DSO performs the optimal power flow calculation based on Eq. 24-31 to determine the flexibility volume $V_{x,t}^{\uparrow},V_{x,t}^{\downarrow}$ to resolve congestion. If the optimization calculation yields a feasible solution, this indicates that the available flexibility from the aggregators is sufficient to resolve the congestion caused by the TSO’s service. Otherwise, the boundary for the TSO is divided by $step$-th in 32 and 33. TSO can only use resources within the new constraints. The process is then repeated until the DSO finds the optimal solution or reaches the stopping condition after five divisions, the $E_{m,t}^{a,\uparrow^{\prime}},E_{k,t}^{a,\downarrow^{\prime}}$ are reset to 0.
@@
Differently, in the DSO-managed coordination scheme, DSO updates the demand and generation at each node with $\overline{E_{x,t}^{a,\uparrow}},\underline{E_{x,t}^{a,\downarrow}}$ from the aggregator. Active power balance condition at node $x$ at step $t$ is presented in Eq. 24. $J_{x}$ refers to the subset of nodes that are connected to node $x$. This constraint needs to be satisfied for all nodes in the network. Similar to the TSO model, the constraints on flexibility boundaries are presented in Eq. 25, Eq. 26, and Eq. 27. Power flow symmetry condition in Eq. 28 indicates that the flow of power from node $x$ to node $y$ ($Flow_{xy}$) is opposite to the value of flow from node $y$ to node $x$. This value is determined by the base power for the network (in MVA) $Base_{MVA}$, the susceptance of line in per unit $B_{xy}$ and the voltage phase angle difference $\theta_{x,t}$ and $\theta_{y,t}$. $B_{xy}$ is calculated by resistance and reactance of the line in Eq. 31. Moreover, the rated value of each branch ($Rated_{xy}$) in Eq. 29 multiplied by the power factor $PF$ forms the thermal rating threshold for each branch. The power factor $PF$ is considered 98% in this study. The optimization problems are mixed-integer linear programming (MILP) and are solved by the Gurobi solver in Pyomo.
@@ IV Case studies
The proposed coordination schemes are validated using the real distribution network at Bleiswijk station in the northern Netherlands. This network comprises 4 step-down two-winding transformers from 25kV to 10.5kV, 182 busbars, 867 terminals, 224 lines, 25 synchronous machines, 170 substations, and 166 loads. The total installed capacity is 66.06 MW, with a total spinning reserve is 11.36 MW. The PV penetration rate is 12.85% (8.5MW) from 16 installation points. The aggregators providing upward regulation service are connected to nodes 12, 42, 145, 146, and 147. While the aggregators providing downward regulation service are connected to nodes 15, 18, 41, 179, and 183 in the network. Each aggregator manages 100 EVs and offers the same total volume of flexibility. The only difference lies in their bid prices, which are presented in Table I.
TABLE I: The DER flexibility for balancing service
@@ IV-A Aggregator behavior by market price and season
One week in summer 2023 (June 26th to July 2nd) and one week in winter 2024 (January 1st to January 7th) are selected for the analysis. With a BRP price of 30 €/MW and a consumer price of 85 €/MW, respectively, which is lower and higher than the average day-ahead price this week. Both in winter and summer, the occurrence of downward service by downward price is very low and has the lowest total volume, followed by upward service, and the highest is the volume purchased from the DA market, which is presented in Fig. 2 and Fig. 3. The distribution of purchase volume by DA price and downward regulation volume by downward price (three figures in the first row) is broader in summer compared to winter (three figures in the second row), primarily due to energy surplus from solar power plants. These surpluses, characterized by zero-marginal prices, lead to many negative price cases for DA and downward price. Contrarily, it’s wider for the distribution of upward regulation service by the upward price range caused by the energy shortage in winter.
Figure 2: Distribution of volume by price in summer and winter.
In winter, although the price upward regulation is high, the decrease in demand for downward service leads to a decrease in upward service, which is demonstrated in Fig. 3. Additionally, when BRP prices are high, downward volume (green part) and upward volume (blue part) significantly decrease, while buy volume (orange part) increases, especially in winter.
Figure 3: Distribution of volume by season and BRP price.
@@
BRP price impact is also illustrated in the total volume upward, downward regulation, and buy from the DA market of each EV. Fig. 4 on the top and bottom shows the different behavior by time (PTU) of an individual EV when the BRP price increases from 30 €/MWh to 200 €/MWh. The orange bars presenting the downward regulation service almost disappear. Additionally, the volume for upward (blue bar) is also decreased, leading to the volume bought from the DA market (green bar) being reduced, because less energy needs to be procured to compensate for the volume discharge when providing upward service.
Figure 4: Volume by type of EV1 with BRP price = 30€/MWh (on the top) and price = 200€/MWh (on the bottom).
IV-B Grid constraint evaluation via two coordination schemes
@@
Providing services brings more benefits to aggregators and prosumers. However, a problem arises when the TSO activates the service without considering the distribution network constraints, and multiple aggregators react at the same time with very high or very low attractive prices, resulting in congestion within the distribution network. The green line in Fig. 5 shows that activating the TSO balancing service at ISP 45 causes overloading in the main line straight to the connection point with the external grid. The congestion above is due to the regulation upward from TSO, the line loading is over 95% due to the upward regulation when the distribution grid is in a low-demand situation, causing a reversed flow from the distribution grid to the transmission grid. This will cause rapid deterioration of the device or damage if the overload condition is prolonged. The dashed red line presents the limitation of the line loading, which is 95% in this study. With different coordination between TSO and DSO, these issues are addressed through different sequences.
In the hybrid scheme, TSO performs an initial filter to select bids before forwarding them to DSO under the MOL form. After DSO validation, the DSO returns updated limits on the volume of flexibility that TSO can use. An example of the MOL at ISP 45 is presented in Fig. 6a. The flexibility boundaries before and after the validation, from ISP 40 to 50, are depicted in Fig. 6b by the blue and red lines, respectively. Loading at the observed line after TSO use flexibility with the new boundaries remains below the constraints threshold, which is presented by the orange line in Fig. 5.
@@
Figure 6: a) MOL upward, b) Flexibility boundaries
Differently, in the DSO-managed model, the aggregator directly sends their flexibility offers to the DSO for validation. DSO performs power flow analysis with the flexibility volume. Then the updated limits are sent to TSO, the simulation results are presented in Fig. 7. Similar to the previous case, before the validation, when DSO updates the forecasted demand with the aggregator offers, there is overloading occurs at ISP 45. Subsequently, the DSO performs a linearized AC optimal power flow (ACOPF) to determine the amount of flexibility required to mitigate congestion, iteratively reducing the flexibility boundaries until a feasible solution is obtained. These updated boundaries are then sent to the TSO. The TSO dispatches bids based on system needs while adhering to the new flexibility limits. The resulting loading at the monitored line after TSO activation is shown by the orange line, which confirms that the grid constraint is now satisfied.
However, the difference between the two approaches is not only the order of validation and dispatch actions, but also the amount of flexibility that the aggregator can provide. In the second case, after DSO validation, the loading is lower than in the previous case. This is because the DSO does not know exactly where the TSO will call services, reduces the boundary across all flexibility units until the congestion can be resolved by the available flexibility. This approach leads to the loss of profit for aggregators if more reserve energy can not be used.
A cost analysis was conducted for both the TSO and the aggregator across different days, are highlighted in Table II. Results show that the hybrid model offers greater benefits for both TSO and the aggregator. Since the DER offer price is cheap, the more reserve energy is used from DER flexibility, the lower the cost to the TSO. Therefore, the hybrid model helps TSO reduce costs by accessing a larger volume of DER flexibility than the DSO-managed model. Furthermore, while more flexibility is provided, the aggregator earns more profit, which explains the higher benefit in the hybrid model compared with the DSO-managed model. Although the reduction in TSO costs may not be significant, the increase in aggregate benefits, ranging from 2% to 11%, is worth considered. Besides the economic benefits, the hybrid scheme requires more complex decision process between TSO and DSO, which may lead to lower DER dispatch efficiency. Furthermore, more information needs to be shared between TSO and DSO, which might require more investment in cybersecurity management.
@@
TABLE II: TSO Cost and Aggregator Benefit under Different coordination schemes
TSO cost (€) Aggregator benefit (€)
Date Hybrid DSO-managed Hybrid DSO-managed
05/May 1.353.194 1.354.360 ($\uparrow 0.1\%$) 15.953 14.786 ($\downarrow 7.3\%$)
18/Jul 671.354 671.681 ($\uparrow 0.1\%$) 17.614 17.284 ($\downarrow 1.9\%$)
14/Aug 68.642 68.908 ($\uparrow 0.4\%$) 2.614 2.349 ($\downarrow 10.1\%$)
05/Sep 690.873 692.622 ($\uparrow 0.3\%$) 15.942 14.187 ($\downarrow 11.0\%$)
06/Sep 614.466 615.237 ($\uparrow 0.1\%$) 19.435 18.661 ($\downarrow 4.0\%$)
@@ V Conclusion
To effectively harness the potential of the DERs’ services in the current context, coordination between TSOs and DSOs is necessary to maximize social welfare as well as minimize the risks of conflicting use cases. Notably, DSOs play a more important role in validating services before being utilized by TSOs. This validation ensures that grid constraints are taken into account when providing services from aggregators under the DSO’s management. Thereby, mitigating the potential risks of unexpected overload caused by the simultaneous reaction of aggregators with prices and requests from TSOs. Analyzing the service provision from the aggregator shows that similar behavior from multiple aggregators in neighboring regions in response to the attractive very high upward regulation price or very low downward regulation price will lead to excessive reversed power flow from their connection points to the higher voltage grid or vice versa. Besides, the price the aggregator contracts with BRP for their real-time deviation has a great impact on the decision to provide services and the final profit. The reduction of downward service demand in winter results in the reduction of upward service and volume purchase from the DA market to compensate for the discharge volume. The coordination between TSO and DSO in the offering process allows the provision of services within the grid constraints. Within the scope of this study, the cost-benefit analysis shows that the hybrid-managed scheme gives an advantage with more volume provided and more benefit to the aggregator and lower cost to the TSO than the DSO-managed scheme. The risks associated with inaccurate forecasts of demand, price, and EV sessions significantly influence the decisions of each participant will be addressed in future studies.